OpenStudy (anonymous):

Which expression is prime? A) x^2 - 9 B) 1 - 9x^2 C) 2a^2 - 25 D) 4x^2 - 9

OpenStudy (cwrw238):

Hint look for expressions of the type a^2 - b^2 They can be factored

OpenStudy (cwrw238):

eg 9 - x^2 = (3-x)(3+x) so thats not prime

OpenStudy (cwrw238):

no 1 - 9x^2 = ( 1- 3x)(1 + 3x)

OpenStudy (cwrw238):

which of the functions does not contain 2 perfect squares?

OpenStudy (cwrw238):

there's just one term not a perfect square

OpenStudy (anonymous):

umm.... c

OpenStudy (cwrw238):

right - 2a^2 is not a perfect square C is prime

OpenStudy (anonymous):

thanks!

OpenStudy (cwrw238):

its not the same type as the other 3 ( i e differnce of 2 squares) and there is no common factor for 2a^2 and 25.

OpenStudy (anonymous):

http://usatestprep.com/modules/gallery/files/5/508/508.gif and is this AAS?